OpenStudy (anonymous):

If f(x) = mx +c and m ≠ 0, find the equation of the corresponding function g(x). Explain why we need the restriction ≠0 in this question. State why a function g(x) does or does not exist in the case m = 0. please explain every step and in detail. thank you very much!

OpenStudy (unklerhaukus):

how is g(x) supposed to be related to f(x) ?

OpenStudy (anonymous):

g(x) is the inverse i think, sorry not sure.

OpenStudy (unklerhaukus):

that would make sense..

OpenStudy (unklerhaukus):

If f(x) = m*x+c and g(x) = f^-1 (x) then x = m*g(x)+c. Solve this for g(x)

OpenStudy (anonymous):

ok,so g(x)=(x-c)/m is that right? how would i explain why we need the restriction ≠0 in this question. State why a function g(x) does or does not exist in the case m = 0.

OpenStudy (anonymous):

i dont understand that last bit

OpenStudy (unklerhaukus):

can you ever divide by zero?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

well, that is why m cannot be zero,

OpenStudy (unklerhaukus):

g(x) = (x-c)/m g(x) is only defined when (x-c)/m is defined, (x-c)/m is not defined when m = 0